what is 5/6 divided by 1/3 can you please help me and show me step by step how to do it
Found 2 solutions by Alan3354, rapture:
Answer by Alan3354(69443)

what is 5/6 divided by 1/3
-----------
Invert the divisor (1/3 for this one), and multiply.
--> (5/6)*3
= 5/2

Answer by rapture(86)

5/6 ÷ 1/3
1. Change the sign from ÷ to x
5/6 × 1/3

2. Invert 1/3 to be 3/1

5/6 x 3/1

3. Multiply numerators and denominators

(5 x 3)/(6 x 1)

15/6
4. Reduce the fraction. This is done by dividing the numerator and denominator by the biggest factor. What is the biggest number that EVENLY DIVIDES into 15 and 6?
How about 3?
15 ÷ 3 = 5
6 ÷ 3 = 2

Answer: 5/2